A car is 4 meter long ,crosses 20 meter long bridge in1.6 seconds .at what constant speed the car was traveling.
1
Expert's answer
2018-01-18T15:50:06-0500
If a car is 4 meter long and crosses 20 meter long bridge in 1.6 seconds, then the constant speed of the car is v = s/t = 20/1.6 = 12.5 meter per second.